OpenStudy (anonymous):

Solve for x. 3(2x - 3) + 4(2 - x) = 0

OpenStudy (anonymous):

Help ?

OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)?

OpenStudy (souvik):

first question..: get all the x terms in one side and the numerical terms in other side

OpenStudy (anonymous):

how?

OpenStudy (souvik):

=3*2x-3*3+4*2-4*x

OpenStudy (anonymous):

how do i solve that?

OpenStudy (anonymous):

-17/2?

OpenStudy (anonymous):

\[17/10\]

OpenStudy (souvik):

6x-9+8-4x=0 now get all the x terms in one side and the numerical terms in other side

OpenStudy (anonymous):

or 1/2

OpenStudy (anonymous):

2x=1 so 1/2

OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)?

OpenStudy (souvik):

right..!

OpenStudy (souvik):

\[\frac{ y-y _{1} }{ x-x _{1} }=\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]